OpenStudy (anonymous):

Which equation does not have the same solution as the others? = 2 x - 9 = 17 x + 12 = 20 2x = 16

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

none of them

OpenStudy (anonymous):

ya it is

OpenStudy (anonymous):

california, 20 - 12 is 8.

OpenStudy (anonymous):

so the first one is different.

OpenStudy (anonymous):

yes ty,but still none of them

OpenStudy (anonymous):

x equals 8 in the second and third equations, but not in the first one. wouldn't that make the first one the odd man out?

OpenStudy (anonymous):

x=17+9=26 x=20-12=8 x=16/2=8

OpenStudy (anonymous):

and he asked which one had a different solution than the others.

OpenStudy (anonymous):

x=6 x=8 x=8

OpenStudy (anonymous):

first had a different solution than second and thrid ty hickninja,I miss understand I think he said which one equal =2

OpenStudy (anonymous):

yeah, not sure what the "=2" is.

OpenStudy (anonymous):

yes that why I miss understand =2